The ratio of length and breadth of a rectangle is 9 : 4 and the breadth is 50 cm less than length. Find the area of the rectangle.
Answer by TimothyLamb(4379)

L/w = 9/4
4L = 9w
w = L - 50
---
4L = 9w
4L = 9(L - 50)
4L = 9L - 50*9
5L = 50*9
L = 10*9
L = 90 cm
w = 40 cm
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---